# 2024_subroc.bib

### v1

@@ 4.3 Discovering Injected Subgroups
Figure 2: Intersection over Union (IoU) of an injected subgroup with subgroups in a top-10 ROC AUC result set, ordered decreasingly by interestingness. “baseline” was obtained searching in a base SubROC setting. The other results were obtained using the full SubROC framework with corresponding cover size weighting $\alpha$ and class balance weighting $\beta$, successfully discovering the injected subgroup. The full SubROC framework accurately found the injected subgroup for any weighting configuration by returning two highly overlapping subgroups in the highest ranks of all result sets.
@@ 4.4 Importance of Cover Size and Class Imbalance Weighting
To illustrate the effect of the interaction between model performance, size and class ratio on the proposed scoring functions, we conduct synthetic experiments (cf. [cook_when_2020]). Data was generated from a bivariate normal distribution with means 00 and variances $1$. The first variable $z$ was used to derive a ground truth label, the second $\hat{z}$ represents the prediction values. Varying their correlation thus varies the “prediction performance” (the $y$-axes). Dataset sizes (cover of a subgroup) are varied in Figures 3(a), 3(b), 3(e), 3(f), 3(i) and 3(j) ($x$ axis) and set to $100$ in Figures 3(c), 3(d), 3(g), 3(h), 3(k) and 3(l). Class balance (NCR) $q$ was varied by thresholding $z$ at the first $q$-quantile in Figures 3(c), 3(d), 3(g), 3(h), 3(k) and 3(l) ($x$ axis) and fixed at $q=0.5$ for Figures 3(a), 3(b), 3(e), 3(f), 3(i) and 3(j). Since scoring functions are relative to the overall dataset, we equivalently derive a dataset with correlation 00, size $100$ and $q=0.5$. The generation process was repeated $20$ times followed by an averaging of scores to reduce noise. Note that this procedure allows to evaluate practically impossible subgroups. When the cover size or NCR of a subgroup is close to that of the overall dataset, then the exceptionality (i.e. performance difference) is limited in practice.
@@ 4.4.1 Skew from Cover Size
Figure 3(a) shows the dependence of $\varphi^{rasl}$ on the number of instances in its input, yielding larger values for larger inputs even if predictive performance is fixed constant (see the horizontal gradient at fixed correlations). This can lead to low values of $\varphi^{rasl}$ on small subgroups, which do not necessarily correspond to a performance improvement over the complete dataset. Underperforming subgroups can therefore not be identified by a fixed threshold on the metric (e.g., $\geq 0$). This means that inherently smaller subgroups have to have a higher error rate to be considered important for ARL, and thus the approach returns results with a higher rate of passing the statistical test (Table 1). However, the strength of this effect can not be influenced by the user.
Neither $\varphi^{rPRAUC}$ nor $\varphi^{rROCAUC}$ change with cover size (3(e), 3(i)). Furthermore these figures show increased noise on the lower end of the cover size axes, indicating increased random effects on these scoring functions for subgroups with few instances which underlines the necessity of the statistical tests introduced by SubROC. Additionally, Figures 3(f) and 3(j) show that the introduced cover size weighting successfully leads to a dependence on input size, encoding a preference for larger inputs for both of these scoring functions.
4.4.2 Skew from Class Ratio
Figures 3(c) and 3(k) show that the ARL as well as the PR AUC depend on the class ratio. The PR AUC skew is known to be related to the unachievable area in PR space [boyd_unachievable_2012]. This skew has a similar influence on the interpretation of the scores as has the cover size skew of the ARL. Figure 3(g) shows a more subtle effect for the ROC AUC through its contour lines, that was also found by [cook_when_2020]. Figures 3(d), 3(h) and 3(l) show that the class balance weighting successfully introduces a preference towards subgroups with balanced classes. In particular, the score of subgroups with highly imbalanced classes is close to zero. Because cover size also has only a minor effect on the ROC AUC (cf. Figure 3(e)), searches utilizing the ROC AUC within SubROC can be expected to yield the most consistently underperforming subgroups over the widest range of cover sizes and class balances compared to the ARL and the PR AUC.
@@ 4.4.2 Skew from Class Ratio
(c) $\varphi^{rasl}_{0,0}$
(d) $\varphi^{rasl}_{0,1}$
(e) $\varphi^{rROCAUC}_{0,0}$
(f) $\varphi^{rROCAUC}_{1,0}$
(g) $\varphi^{rROCAUC}_{0,0}$
(h) $\varphi^{rROCAUC}_{0,1}$
(i) $\varphi^{rPRAUC}_{0,0}$
(j) $\varphi^{rPRAUC}_{1,0}$
(k) $\varphi^{rPRAUC}_{0,0}$
(l) $\varphi^{rPRAUC}_{0,1}$
Figure 3: Skewed response of cover size and class balance weighted scoring functions depending on cover size, NCR and correlation (model performance) on synthetic data. Contour lines show smoothed paths along which interestingness is equal.
@@ 5 Related Work
[sagadeeva_sliceline_2021] only consider residuals-based loss functions to rank subgroups in contrast to our ranking-based measures. Similarly, [zhang_identifying_2017] does not use AUC metrics or the subgroup discovery framework. Other research allows for user-specified performance measures for subgroup ranking, but they do not provide specific support such as optimistic estimates or generalization-awareness[chung_automated_2020, zhang_sliceteller_2017]. [cabrera_fairvis_2019, chung_automated_2020, zhang_sliceteller_2017] propose automatic search approaches as part of generic interactive evaluation tools. None of these works applied the difference between the subgroup AUC to the overall dataset AUC for the ROC or PR curve in a subgroup discovery setting and investigated its properties in detail.
The following works are concerned with the ROC curve of subgroups in detail. [gardner_evaluating_2019] propose to use the area between the ROC curves of a model on two subgroups in the context of fairness of data slicing but do not discuss the mining of respective patterns. [carrington_deep_2023] propose to evaluate specific parts of the ROC curve and call these parts subgroups. Considered parts in that work are intervals in the x-axis or the y-axis of the ROC space as well as intervals in the thresholds used to compute the ROC points from. In addition to this differing notion of subgroups, they do not investigate an automated search for subgroups. [tafvizi_attributing_2022] presents methods for computing the attribution of subgroups to the overall AUC. They do this to identify parts of the data that contribute to a reduction in the overall AUC. The contribution of a subgroup to the overall dataset AUC is highly influenced by how the subgroup is embedded into the ordering of the overall dataset. Therefore finding parts of the dataset with an exceptional AUC would simply not contribute to their goal. Our goal is to find under which circumstances a model is most likely to fail or succeed, so we take a different approach with AUC-based exceptionality.
For subgroup discovery, many (tight) optimistic estimates (OE) have been proposed, primarily for settings with traditional target concepts (binary, numeric, or both) [grosskreutz_fast_2010, lemmerich_fast_2016]. In a few settings, tight OE have also been derived for more complex model classes in the EMM framework [kalofolias_discovering_2019, belfodil_identifying_2020]. An OE for a scoring function that computes the ROC AUC is given in [lemmerich_fast_2016]. However, there the ROC curve is based on different attributes than what is studied in this work, such that the OE is not applicable here. To the best of our knowledge, none of the previously presented OEs can be applied to the scoring functions used in this work.
@@ 6 Conclusion
We introduced SubROC, a novel framework for reliable and efficient discovery of interpretable subgroups with exceptional model performance with respect to soft classifiers. Our experiments demonstrate the necessity for such a framework as well as the effectiveness of SubROC when using the performance measures ARL, ROC AUC and PR AUC. We identified the side effects of a subgroup’s cover size and class balance on these performance measures and their intricate influence on the results. By providing optimistic estimates, we enable fast mining. To summarize, SubROC has the potential to enable routine evaluation of ML models, improve quality control procedures, inform model performance enhancements, and increase confidence in deployed ML models overall.
In this work, we focused on ranking-based soft classifier evaluation measures due to their widespread use and their specific challenges to Exceptional Model Mining. Other popular performance measures, such as cross-entropy or the Brier score, can be integrated into classic subgroup discovery. Their in-depth evaluation is an avenue for future research. Also, directly integrating statistical tests into the search may help to further improve the significance of the discovered subgroups. Finally, the investigation of alternatives to account for class imbalances is a promising direction for future work.
@@ 7 Acknowledgements
The funding for this study was in part provided by the BMBF (01IS22077).
[…]
@@ Appendix A Proofs of Tight Optimistic Estimates
Our formalizations make use of standard multiset notation as presented in [syropoulos_mathematics_2001]. To handle values of instances in a dataset $D=(\mathcal{I},\mathcal{A})$ we define a database as the multiset $d(D)=[(A_{1}(c),\dots,A_{|\mathcal{A}|}(c))|c\in\mathcal{I}]$. A shorthand notation for ground truth and prediction values is $I^{\{y,\hat{y}\}}=[(Y(c),\hat{Y}(c))|c\in I]$ on a set of instances $I$, the item values of which are referenced by $y_{i}$ and $\hat{y}_{i}$ with index $i$.
@@
The main interestingness measure $\varphi^{rasl}$ from [duivesteijn_understanding_2014] is defined with alternative signatures (compared to the main text) for use in the proofs as
@@
$\varphi^{rasl}(p):=$ $ARL(sg(p)^{\{y,\hat{y}\}})-ARL(\mathcal{I}^{\{y,\hat{y}\}})$
for multisets of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$. $\varphi^{rasl}$ is a special case of Definition 3 with $m_{l}=ARL$. The definitions from the main text can be formalized on this basis as
$PEN(c,p):=$ $PEN_{index(c)}(sg(p)^{\{y,\hat{y}\}})$
$ARL(p):=$ $ARL(sg(p)^{\{y,\hat{y}\}})$
@@
$\varphi^{ROCAUC}$ and $\varphi^{PRAUC}$ are both instances of Definition 2 with the ROC AUC and the PR AUC performance measures substituted in place of $m_{s}$ respectively. $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ are the same for Definition 4.
@@
Definition 5 (ROCAUC [fawcett_introduction_2006]).
@@
Definitions of (tight) lower bounds of soft classifier performance measures are now given. They allow to structure the construction of some of the tight optimistic estimates that are presented in this section into a statement about a performance measure, considering a bound for the performance measure, and an adaptation step of that bound to interestingness measures, yielding an optimistic estimate. The lower bound definition (Definition 8) requires for a soft classifier performance measure an additional soft classifier performance measure that does not return smaller values than the reference performance measure for any input. The tightness criterion of Definition 9 requires a lower bound to be equal to the reference performance measure for at least one input. The performance measure lower bound definition and the bound tightness definition are defined analogously to optimistic estimates and tight optimistic estimates respectively.
@@
[lemmerich_novel_2014] states that for all two rules $p_{gen}$ and $p_{spec}$ the following statement holds
@@
Equation 3 follows from Equation 4 and Equation 5, concluding the proof for Equation 1.
Proof for Equation 2
@@
$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge m_{l}(sg(p^{\prime})^{\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$
@@
From Equation 7 and Equation 8 follows
$\exists C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}\exists\mathcal{I}^{\prime}\subseteq sg(p):C^{\prime}=\mathcal{I^{\prime}}^{\{y,\hat{y}\}}\wedge m_{l}(C^{\prime})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (9)
The indifference of the order in a series of the same quantifiers, the equality of $C^{\prime}$ and $\mathcal{I^{\prime}}^{\{y,\hat{y}\}}$ as well as simplification rules imply that the following is equivalent to Equation 9
$\exists\mathcal{I}^{\prime}\subseteq sg(p):m_{l}(\mathcal{I^{\prime}}^{\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$
This is trivially equivalent to Equation 6, concluding the proof for Equation 2. ∎
@@
It follows from Lemma 7 that the soft classifier loss interestingness measure $\varphi^{m_{s}^{\prime}}$ based on $m_{s}^{\prime}$ has the following tight optimistic estimate
$oe_{\varphi^{m_{s}^{\prime}}}(p)=b_{m_{s}^{\prime}}(sg(p)^{\{y,\hat{y}\}})$
@@
Equation 12 is true if and only if Equation 13 is true, concluding the proof for Equation 10.
Proof for Equation 11
@@
Equation 14 is true if and only if Equation 15 is true, concluding the proof for Equation 11. ∎
All in all by sequentially applying Lemma 7 (respectively Lemma 8) and then Lemma 9 to a soft classifier performance measure upper (respectively lower) bound for a performance measure $m$ we obtain a tight optimistic estimate for the relative soft classifier loss (respectively score) interestingness measure based on $m$. This proving scheme is used to prove tight optimistic estimates for the relative ARL, ROC AUC and PR AUC based interestingness measures.

### v2

@@ 4.3 Discovering an Injected Subgroup
Figure 2: Intersection over Union (IoU) of subgroup covers, comparing an injected subgroup with subgroups in a top-10 ROC AUC result set, ordered decreasingly by interestingness. “baseline” was obtained searching in a base SubROC setting. The other results were obtained using the full SubROC framework with corresponding cover size weighting $\alpha$ and class balance weighting $\beta$, successfully discovering the injected subgroup. The full SubROC framework accurately found the injected subgroup for any weighting configuration by returning two highly overlapping subgroups in the highest ranks of all result sets.
To investigate the validity of SubROC and its ability to discover expected subgroups, we injected an artificially underperforming subgroup into the Adult dataset based on the predictions of an xgboost classifier. The subgroup was randomly drawn from candidates with pattern length $\leq 3$ whose cover size fell between $0.4\%$$0.6\%$ of the search set. It covers 54 instances, has an NCR of 0.2, and an ARL of 1.07, PR AUC of 0.97, and ROC AUC of 0.90 before and 9.93, 0.62, and 0.1, respectively, after injection. For the injection, we multiplied the predicted scores in the subgroup by 1-1. Thereby, the injected subgroup forms a known weakness in the predictions, and we can evaluate the subgroup search by whether this subgroup is included in the result set. The subgroup discovery was performed up to depth 4 under different configurations (baseline without the SubROC framework and three different weightings of cover size and class balance in the SubROC framework). Result sets were top-10 filtered to reflect interpretable result set sizes.
Figure˜2 shows the IoU as an overlap measure between found subgroups and the injected subgroup for the ROC AUC. In contrast to the baseline, the SubROC framework successfully discovers the injected subgroup and reports it at the highest ranks in the result set. Note that multiple patterns can describe the same set of instances, so the injected errors can be found several times.
While ROC AUC robustly identifies the subgroup, similar experiments on ARL or the PR AUC did not reproduce these findings (all returned $\text{IoU}<0.1$; see Section˜B.4). The injected subgroup contains mostly negatives, which interacts with the class balance skews of ARL and PR AUC (see Section˜4.4.2). Furthermore, the injected subgroup is small, making it inherently hard to capture with ARL, which has an embedded size skew (see Section˜4.4.1). This highlights the importance of the size and class balance terms introduced in SubROC to counter these skews, but also shows that they do not completely remove the effects. Future work, for example, employing a normalized PR AUC [16] may provide further improvement.
@@ 4.4 Importance of Cover Size and Class Imbalance Weighting
To illustrate the effect of the interaction between model performance, size and class ratio on the proposed scoring functions, we conduct experiments on synthetic data (cf. [15]). Data was generated from a bivariate normal distribution with means 0 and variances $1$. The first variable $z$ was used to derive a ground truth label, the second $\hat{z}$ represents the prediction values. Varying their correlation thus varies the “prediction performance” (the $y$-axes). Dataset sizes (cover of a subgroup) are varied in Figure˜3 ($x$ axis) and set to $100$ in Figure˜4. Class balance (NCR) $q$ was varied by thresholding $z$ at the first $q$-quantile in Figure˜4 ($x$ axis) and fixed at $q=0.5$ for Figure˜3. Since scoring functions are relative to the overall dataset, we equivalently derive a dataset with correlation 0, size $100$ and $q=0.5$. The generation process was repeated $20$ times followed by an averaging of scores to reduce noise. Note that this procedure allows to evaluate practically impossible subgroups, since any subgroup with cover size or NCR close to that of the overall dataset can only have limited exceptionality (i.e. performance difference) in practice.
@@ 4.4.1 Skew from Cover Size
Figure˜3(a) shows the dependence of $\varphi^{rasl}$ on the number of instances in its input, yielding larger values for larger inputs even if predictive performance is kept constant (see the horizontal gradient at fixed correlations). This can lead to low values of $\varphi^{rasl}$ on small subgroups, which do not necessarily correspond to a performance improvement over the complete dataset. Underperforming subgroups therefore cannot be identified by a fixed threshold on the metric (e.g., $\geq 0$). Therefore smaller subgroups have to have a higher error rate to be considered important for ARL, and thus the approach returns results with a higher rate of passing the statistical test (Table˜1). However, the strength of this effect can not be controlled by the user.
Neither $\varphi^{rPRAUC}$ nor $\varphi^{rROCAUC}$ change with cover size (3(c), 3(e)). Furthermore these figures show increased noise on the lower end of the cover size axes, indicating increased random effects on these scoring functions for subgroups with few instances which underlines the necessity of the statistical tests introduced by SubROC. Additionally, Figures˜3(d) and 3(f) show that the introduced cover size weighting leads to a dependence on input size, encoding a preference for larger inputs for both of these scoring functions.
@@
(c) $\varphi^{rROCAUC}_{0,0}$
(d) $\varphi^{rROCAUC}_{1,0}$
(e) $\varphi^{rPRAUC}_{0,0}$
(f) $\varphi^{rPRAUC}_{1,0}$
Figure 3: Skewed response of cover size and class balance weighted scoring functions depending on cover size and correlation (model performance) on synthetic data. Contour lines show smoothed paths along which interestingness is equal. Unweighted scoring functions based on the AUC performance measures show increased noise on subgroups with small covers. Significant dependencies on the cover size are visible in the results for the unweighted scoring function based on the ARL performance measure as well as all weighted scoring functions.
(a) $\varphi^{rasl}_{0,0}$
(b) $\varphi^{rasl}_{0,1}$
(c) $\varphi^{rROCAUC}_{0,0}$
(d) $\varphi^{rROCAUC}_{0,1}$
(e) $\varphi^{rPRAUC}_{0,0}$
(f) $\varphi^{rPRAUC}_{0,1}$
Figure 4: Skewed response of cover size and class balance weighted scoring functions depending on negative class ratio and correlation (model performance) on synthetic data. Contour lines show smoothed paths along which interestingness is equal. The unweighted scoring functions based on the ARL and the PR AUC both increase with the NCR for almost any correlation. The unweighted scoring function based on the ROC AUC depends on the difference between the NCR and $0.5$. All weighted scoring functions return the most extreme values close to NCR s of $0.5$, with a gradient from low to high scores as correlations decrease.
4.4.2 Skew from Class Ratio
Figures˜4(a) and 4(e) show that the ARL and PR AUC depend on the class ratio. The PR AUC skew is known to be related to the unachievable area in PR space [16]. This skew has a similar influence on the interpretation of the scores as has the cover size skew of the ARL. Figure˜4(c) shows a more subtle effect for the ROC AUC through its contour lines, that was also found by [15]. Figures˜4(b), 4(d) and 4(f) show that the class balance weighting introduces a preference towards subgroups with balanced classes. In particular, the score of subgroups with highly imbalanced classes is close to zero. Since ROC AUC exhibits the least prominent skews (cf. Figures˜3(c) and 4(c)), searches utilizing the ROC AUC within SubROC can be expected to yield the most consistently underperforming subgroups over the widest range of cover sizes and class balances compared to the ARL and PR AUC.
@@ 5 Related Work
Sagadeeva and Boehm [30] only consider residuals-based loss functions to rank subgroups in contrast to our ranking-based measures. Similarly, Zhang and Neill [31] use neither AUC metrics nor the subgroup discovery framework. Other research allows for user-specified performance measures for subgroup ranking, but they do not provide specific support such as optimistic estimates or generalization-awareness[6, 32]. [7, 6, 32] propose automatic search approaches as part of generic interactive evaluation tools. None of these works applied the difference between the subgroup AUC to the overall dataset AUC for the ROC or PR curve in a subgroup discovery setting and investigated its properties in detail.
The following works are concerned with the ROC curve of subgroups in detail. [33] propose to use the area between the ROC curves of a model on two subgroups in the context of fairness of data slicing but do not discuss the mining of respective patterns. [34] propose to evaluate specific parts of the ROC curve and call these parts subgroups. Considered parts in that work are intervals in the x-axis or the y-axis of the ROC space as well as intervals in the thresholds used to compute the ROC points from. In addition to this differing notion of subgroups, they do not investigate an automated search for subgroups. [35] presents methods for computing the attribution of subgroups to the overall AUC. They do this to identify parts of the data that contribute to a reduction in the overall AUC. The contribution of a subgroup to the overall dataset AUC is highly influenced by where its instances fall within the global ranking of predictions for the entire dataset. Therefore finding parts of the dataset with an exceptional AUC would simply not contribute to their goal. Our goal is to find under which circumstances a model is most likely to fail, so we take a different approach with AUC-based exceptionality.
For subgroup discovery, many (tight) optimistic estimates (OE) have been proposed, primarily for settings with traditional target concepts (binary, numeric, or both) [36, 37]. In a few settings, tight OE have also been derived for more complex model classes in the EMM framework [38, 39]. An OE for a scoring function that computes the ROC AUC is given in [37]. However, there the ROC curve is based on different attributes compared to ours, such that the OE is not applicable here. To the best of our knowledge, no previously presented OE can be applied to the scoring functions used in this work.
@@ 6 Conclusion
We introduced SubROC, a novel framework for reliable and efficient discovery of interpretable subgroups with exceptional model performance for soft classifiers. Our experiments demonstrate the necessity for such a framework as well as the effectiveness of SubROC when using the performance measures ARL, ROC AUC and PR AUC. We identified the side effects of a subgroup’s cover size and class balance on these performance measures and their intricate influence on the results. By providing optimistic estimates, we enable fast mining. To summarize, SubROC has the potential to enable routine evaluation of ML models, improve quality control procedures, inform model performance enhancements, and increase confidence in ML models.
In this work, we focused on ranking-based soft classifier evaluation measures given their widespread use and the specific challenges they pose for integration into the Exceptional Model Mining framework. By contrast, other popular performance measures, such as cross-entropy or the Brier score decompose over individual instances and can therefore be more directly integrated into classic subgroup discovery. Their in-depth evaluation is an avenue for future research. Also, directly integrating statistical tests into the search may help to further improve the significance of the discovered subgroups. Finally, the investigation of alternatives to account for class imbalances is a promising direction for future work.
@@ 7 Acknowledgements
Funding was provided by the BMFTR (01IS22077, 01ED2507).
@@ Appendix A Proofs of Tight Optimistic Estimates
Our formalizations make use of standard multiset notation as presented in [40]. To handle values of instances in a dataset $D=(\mathcal{I},\mathcal{A})$ we define a database as the multiset $d(D)=[(A_{1}(c),\dots,A_{|\mathcal{A}|}(c))|c\in\mathcal{I}]$. A shorthand notation for ground truth and prediction values is $I^{\{y,\hat{y}\}}=[(Y(c),\hat{Y}(c))|c\in I]$ on a set of instances $I$, the item values of which are referenced by $y_{i}$ and $\hat{y}_{i}$ with index $i$.
@@
The main interestingness measure $\varphi^{rasl}$ from [8] is defined with alternative signatures (compared to the main text) for use in the proofs as
@@
[…]
for multisets of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$. $\varphi^{rasl}$ is a special case of Definition˜3 with $m_{l}=ARL$. The definitions from the main text can be formalized on this basis as
[…]
@@
$\varphi^{ROCAUC}$ and $\varphi^{PRAUC}$ are both instances of Definition˜2 with the ROC AUC and the PR AUC performance measures substituted in place of $m_{s}$ respectively. $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ are the same for Definition˜4.
@@
Definition 5 (ROCAUC [41]).
@@
Definitions of (tight) lower bounds of soft classifier performance measures are now given. They allow to structure the construction of some of the tight optimistic estimates that are presented in this section into a statement about a performance measure, considering a bound for the performance measure, and an adaptation step of that bound to interestingness measures, yielding an optimistic estimate. The lower bound definition (Definition˜8) requires for a soft classifier performance measure an additional soft classifier performance measure that does not return smaller values than the reference performance measure for any input. The tightness criterion of Definition˜9 requires a lower bound to be equal to the reference performance measure for at least one input. The performance measure lower bound definition and the bound tightness definition are defined analogously to optimistic estimates and tight optimistic estimates respectively.
@@
[12] states that for all two rules $p_{gen}$ and $p_{spec}$ the following statement holds
@@
Equation˜3 follows from Equation˜4 and Equation˜5, concluding the proof for Equation˜1.
Proof for Equation˜2
@@
[…]
@@
From Equation˜7 and Equation˜8 follows
[…]
The indifference of the order in a series of the same quantifiers, the equality of $C^{\prime}$ and $\mathcal{I^{\prime}}^{\{y,\hat{y}\}}$ as well as simplification rules imply that the following is equivalent to Equation˜9
[…]
This is trivially equivalent to Equation˜6, concluding the proof for Equation˜2. ∎
@@
It follows from Lemma˜7 that the soft classifier loss interestingness measure $\varphi^{m_{s}^{\prime}}$ based on $m_{s}^{\prime}$ has the following tight optimistic estimate
[…]
@@
Equation˜12 is true if and only if Equation˜13 is true, concluding the proof for Equation˜10.
Proof for Equation˜11
@@
Equation˜14 is true if and only if Equation˜15 is true, concluding the proof for Equation˜11. ∎
All in all by sequentially applying Lemma˜7 (respectively Lemma˜8) and then Lemma˜9 to a soft classifier performance measure upper (respectively lower) bound for a performance measure $m$ we obtain a tight optimistic estimate for the relative soft classifier loss (respectively score) interestingness measure based on $m$. This proving scheme is used to prove tight optimistic estimates for the relative ARL, ROC AUC and PR AUC based interestingness measures.